The average salary for computer progammers is $7740 less than twice the average salary for carpenters. Combined, their average salaries are $99000.Determine the averaage salary for each of these job.
Answer by nerdybill(5411)

The average salary for computer progammers is $7740 less than twice the average salary for carpenters. Combined, their average salaries are $99000.Determine the average salary for each of these job.
.
Let x = salary of a carpenter
then
2x-7740 = salary of a computer programmer
.
[x + (2x-7740)]/2 = 99000
[3x-7740]/2 = 99000
3x-7740 = 198000
3x = 205740
x = $68580 (carpenter's salary)
.
To figure out the salary for a computer programmer:
2x-7740
= 2(68580)-7740
= 137160-7740
= $129420 (programmer's salary)